OpenStudy (baby456):

4x+8 - 6= 2 (9-2) help me please for a medal

OpenStudy (tkhunny):

With what are you struggling? Why not do the subtraction in the parentheses, first?

OpenStudy (baby456):

okay

OpenStudy (anonymous):

4x+8-6=2(9-2) 4x+2=2(7) 4x-2=14 4x=16 x=4

OpenStudy (baby456):

thanks

OpenStudy (tkhunny):

Thanks for doing your homework for you? @jewbitz, please read the terms of use for this discussion board.

OpenStudy (baby456):

i told her to help

OpenStudy (tkhunny):

The definition of "Help" on this website is NOT just to do all the work. It is to HELP you do your own work. When it is all done for you, you learn less and that is not satisfactory.

OpenStudy (baby456):

i actually sdid some of it myself just needed to see if it was rightr

OpenStudy (tkhunny):

This is a common claim. Unfortunately, you did not share any of that with us. Please show YOUR work in the beginning. This ALWAYS helps.